Find I, and ly for the composite area shown. Show complete solution with formulae as needed. - 3 in. -10.5 in
Answer #1

the centroids of Solution: At first we need to find composite area. k 3.5 in MY 13.5/2= 6.75in 3 in 340.5 -1.5in IS 10.5 in 4so = I xi (31.5x5-25)+(1575*4.5) ZAD + (40.5x6.75) (31-5+ 15.75 + 40.5) on n = 5.808 in also, - IDEAL (31-5*1-5)+(1575*8.25)Ix For 4 for Calculation: Part 1; A = 31.5 in 2 (1x) = 10.5 x 3 12 doi= (5-4,2 on ay = (8.942-10572 on, da = (7.442)2 in2 PosThere force; x = [ (IX)o, + Ady, + [(Iy). + A, dy 2] + [(1x)og+ Az dy] 09 Ix = 10.5 x 3² + 31.5x (7.442) 2 + 105x10-5} + 1dx2 - (n_n) 2 daz² = (5.808 - 4.52 da 2 2 = (1.30872 in2 For Part 3; Az = 40.5 in 2 (Ig) o = 3x13.53 ist daze = (in – x3)2 da